A population of values has a normal distribution with μ=207.3 and σ=35.8. A random sample of size n=103 is drawn. Find the probability that a sample of size n=103 is randomly selected with a mean less than 208.4. Round your answer to four decimal places.

P(M<208.4)=
Answer #1

Solution :

Given that ,

mean =   = 207.3

standard deviation = =   35.8.

n = 103

= 207.3

=  / n = 35.8/ 103=3.53

P(M <208.4 ) = P[(M - ) / < (208.4-207.3) /3.53 ]

= P(z <0.31 )

Using z table  

= 0.6217   

probability= 0.6217